OpenStudy (anonymous):

need help finding the triple integral 0-1 for each (4x^2y+xyz+yz^2)dxdydz

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{1}\int\limits_{0}^{1}(4x^{2}y+xyz+yz ^{2})dxdydz\]

OpenStudy (anonymous):

Did you try just doing it directly?

OpenStudy (anonymous):

i don't know how to

OpenStudy (anonymous):

\[\int_0^1\left(\int_0^1\left(\color{red}{\int_0^1 \left(4x^2y+xyz+yz^2\right)~dx}\right)~dy\right)~dz\] Take it one variable at a time, treating the other variables as constants.

OpenStudy (anonymous):

By "directly", I mean for instance, you don't do anything to change the order of integration, you just do what siths said.

OpenStudy (anonymous):

I see, I'll give that a try, thanks